How many grams of CO are needed to react with 1.84 mol of Fe2O3? Express the mass in grams to three significant figures.


1
Expert's answer
2022-02-10T20:38:03-0500

Fe2O3 + 3CO → 2Fe + 3CO2

Mole ratio= 1.3

If 1.84 moles of Fe2O3 react then 1.84× 3= 5.52moles of CO react

Mass= moles× rmm

Mass= 5.52× 28= 152.560g